Let (2, F. P) be a probability space and let E, and Ey be two events with P(E₁) = 0.2. P(E₂) 0.4 and P(E₁ E₂)=0.1.


Find the probability that (a) Exactly one of the events E, or E₂ will occur.


(b) At least one of the events E, or E, will occur.


(c) None of E, and Es will occur.


1
Expert's answer
2021-08-17T07:28:15-0400

Solution:

P(E₁) = 0.2. P(E₂) 0.4 and P(E₁"\cap" E₂)=0.1.

(a) Exactly one of the events E, or E₂ will occur.

Required probability = P(E₁) + P(E₂) - 2P(E₁"\cap" E₂)

"=0.2+0.4-2\times0.1
\\=0.6-0.2
\\=0.4"

(b) At least one of the events E1 or E2 will occur.

Required probability = P(E₁"\cup" E₂) = P(E₁) + P(E₂) - P(E₁"\cap" E₂)

"=0.2+0.4-0.1
\\=0.6-0.1
\\=0.5"

(c) None of E1 and E2 will occur.

Required probability = P(E1' "\cap"E2') = P[(E₁"\cup" E₂)'] = 1-P[(E₁"\cup" E₂)]

"=1-0.5
\\=0.5"